Find the optical density and transmittance of 0.04%-lactise solution, if 1/8 of initial light intensity is absorbed in it


1
Expert's answer
2021-04-26T17:13:44-0400

Transmittance:


"T=1-\frac{1}{8}=0.875"

The optical density (or absorbance) is then:


"D=\u2212\\log _{10}\n\u200b\t\n T=\u2212\\log _{10}\n\u200b\t\n 0.875\u22480.058"